Determine the functions sin, cos and tan and their reciprocal and inverse functions.


1
Expert's answer
2021-05-10T13:02:39-0400

"\sin \theta = \frac{{Opposite}}{{Hypothenuse}}\\ 
 \cos\theta = \frac{{Adjacent}}{{Hypothenuse}}\\ 
 \tan\theta = \frac{{Opposite}}{{Adjacent}} = \frac{\sin\theta}{\cos\theta} \\"

Cosecant, secant and tangent are the reciprocals of sine, cosine and tangent.

Arcsine, arccosine and arctangent are the inverse functions of sine, cosine and tangent.